OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches zero of quantity x squared plus three divided by x to the fourth power.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ x^2+3 }{ x^4 }\]

OpenStudy (solomonzelman):

you can take the limit of the numerator, and then limit of the denominator. (I think )

OpenStudy (anonymous):

No, you can't do that because the denominator tends to 0.

OpenStudy (solomonzelman):

I mean this, \[\lim_{x \rightarrow 0 }~\frac{x^2+3}{x^4}\] \[\Large~\frac{ \lim_{x \rightarrow 0 }~(x^2+3)}{\lim_{x \rightarrow 0 }x^4}\] \[\Large~\frac{ \lim_{x \rightarrow 0 }~(x^2+3)}{(\lim_{x \rightarrow 0 }x)^4}\]

OpenStudy (solomonzelman):

and it would be equivalent to ∞

OpenStudy (anonymous):

You can separate them, but it doesn't help solve the problem.

OpenStudy (solomonzelman):

yes, it doesn't.

OpenStudy (solomonzelman):

I don't find anything better though.

OpenStudy (anonymous):

Well, normally what you would do is factor the top in bottom and then cancel out any repeated roots. If the number approached is still a root of the denominator, then you have an asymptote. This means the limit will be undefined, infinite, or negatively infinite.

OpenStudy (solomonzelman):

yes, it is ∞, because it is unfactorable...

OpenStudy (solomonzelman):

(In this case ∞)

OpenStudy (anonymous):

In this case we can use: \[ \lim_{x\to 0}\frac 1{|x|} = \infty \]If we really want to use limit rules.

OpenStudy (anonymous):

\[ \lim_{x\to 0}\frac{x^2+3}{x^4} = \lim_{x\to0}\frac{1}{x^2} +3 \lim_{x\to0}\frac{1}{x^4} = \left(\lim_{x\to0}\frac{1}{|x|} \right)^2+3 \left(\lim_{x\to0}\frac{1}{|x|}\right)^4 = \infty \]

OpenStudy (anonymous):

Though even doing that is a bit sketchy due to the fact that the limits don't approach a finite number.